OpenStudy (anonymous):

1. If I know the lengths of two sides of a right triangle, how do I find the third? 2. Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? 3. Could I find the two missing angle measures if I know some of the side lengths of a right triangle? 4. What makes a triangle a “special” right triangle? How can special right triangles help me find side lengths? (Sorry to those who posted on my last question but the site mysteriously deleted it :[ )

OpenStudy (experimentx):

1) use Pythagoras theorem p^2+b^2 = h^2 2) yes you can 3) if you know two sides, then you can find missing angle 4) what is 'special' right angle triangle? FYI: if you know one side .... and any other (angle or side) ... and you can find anything on right angle triangle. just two parameters

OpenStudy (anonymous):

Can you elaborate on 2 and 3?

OpenStudy (anonymous):

Formulas, Plug-ins (etc).

OpenStudy (experimentx):

I told you ... you have three trigonometric ratios sin(Q) = p/h, cos(Q) = b/h tan(Q) = p/b the other angle would be (90-Q) sin(90-Q) = b/h cos(90-Q) = p/h tan(90-Q) = b/p choose your ratio connecting the side and angle you have known, find the missing other ... and move on to other ratio

OpenStudy (anonymous):

I know you did, but for some reason the question was deleted and I lost all of my replies :\ thank you for the help